• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 16:04:34 on March30,2005

    Entry number 142809

    keyword=Day Shift Summary

    At start of shift everything running smoothly at 30 microamps
    "F1" kinematics
    Pleft=3.724 19.5 deg
    Pright=1.42 35.805 deg
    BB=518.375 amps 0.932 T

    8:25 request 50 microamps. MCC says they can only get 45. We run at 45.
    09:00 Beam off for tests and spot move. MCC estimates 1 hour to complete.
    10:02 MCC says they're ready to restore beam. Will start with a trickle. If all looks OK we go to 50 Microamps.
    10:36 Lost Q1 and dipole on right side. Requires access to reset
    11:30 Folts says everything is back up. Problem was caused by a hiccup at ESR. We just have to wait for the right dipole to go to zero and come back up.
    11:35 Right dipole power supply faults when we try to turn it on. Needs a quick access to reset.
    11:57 ramping up.
    12:00 Riad uses the available beam for LUMI studies. See halog 142802
    We run the DAQ (run 3072)just to collect information
    13:49 Riad is finished and right dipole is back (good timing)
    BUT, need controlled access to reset SRC HV.
    14:25 OK everthing is back up. Except, we have to reboot some roc's to get the DAQ running
    14:38 beam is back and DAQ is running.

    Run summary:
    3069 07:24-08:10
    3070 08:11-09:00
    3071 10:06-10:36 end with spectrometer magnet trips: right dipole and right Q1
    3072 12:27-12:37 Junk Right Dipole ramping

    3074 14:38-15:15
    3075 15:15-15:53
    3076 15:54-